The components of a velocity field are given by u = x + y, and v = xy3 + 81 and w = 0. Determine the location of the stagnation point (V = 0) in the flow field where y is positive
Answer #1

slagnation poort reloaty =o u=0 ivao o=x-ty x=-4 xy & so 1-y. y3_-8) y=3m] x=-Y x=-3ml location (x,y) = (-315) n